### v1

Abstract
We examine the location characteristics of a conditional selective confidence interval based on the polyhedral method. This interval is constructed from the distribution of a test statistic conditional upon the event of statistical significance. In the case of a one-sided test, the behavior of the interval varies depending on whether the parameter is highly significant or only marginally significant. When the parameter is highly significant, the interval is similar to the usual confidence interval derived without considering selection. However, when the parameter is only marginally significant, the interval falls into an extreme range and deviates greatly from the estimated value of the parameter. In contrast, an interval conditional on two-sided significance does not yield extreme results, although it may exclude the estimated parameter value.
Keywords: Selective inference, statistical significance, confidence interval, polyhedral method. Multiple hypothesis testing, post-selection inference, conditional inference.
Lately, selective inference has received increased attention from researchers. A partial list of papers includes Berk et al. (2013); Lockhart et al. (2014); Fithian et al. (2017); Tibshirani et al. (2016); Lee et al. (2016); Bachoc et al. (2019); Watanabe & Suzuki (2021); Jewell et al. (2022); Terada & Shimodaira (2023); Duy & Takeuchi (2023), and a review is available in Zhang et al. (2022).
Methods for selective inference provide valid statistical inference on parameters when the estimated model or the parameters of interest are selected based on the observed data. Conventional inference assumes that the model or parameters are fixed in advance. Empirical practice often diverges from this ideal, potentially invalidating conventional methods. For example, standard confidence intervals for parameters that are chosen based on the observed data may not have the desired coverage. Data-dependence may be introduced by formal methods for model selection such as the Lasso (Lee et al., 2016) or by more informal approaches of data-mining. An example of informal data-mining is a $p$-hacking scenario where a researcher estimates many parameters and then crafts a narrative highlighting the few that are statistically significant.
While various methods have been proposed for selective inference, there has been limited research on the specific properties of the results produced by these methods with notable exceptions by Kivaranovic & Leeb (2021, 2024). This note addresses such a gap in the literature by examining the characteristics of conditional selective confidence intervals constructed using the polyhedral approach described by Lee et al. (2016). We are particularly interested in the behavior of the confidence intervals at the margin of the conditioning set. Our main finding is that the conditional selective confidence interval may be positioned in extreme ranges and can deviate greatly from the conventional parameter estimate.
Our setting is the “file drawer problem” (Rosenthal, 1979), that is, the situation where only statistically significant estimates are presented. The main result is on inference on a parameter $\theta$ that is found significant by a one-sided $t$-test. Let $X$ denote the corresponding $t$-statistic and $\bar{x}$ the critical value. Suppose that we observe $X=x^{\text{obs}}$. The parameter is significant if $x^{\text{obs}}\geq\bar{x}$. The polyhedral method for selective inference on $\theta$ is based on the distribution of $X$ conditional on observing $X\geq\bar{x}$, which has the distribution function $F(a;\theta,\bar{x}):=\Pr(X\leq a\mid X\geq\bar{x})$. Note that $F$ depends on $\theta$ because of the dependence of the distribution of $X$ on it. The conditional confidence interval is found by evaluating the distribution function at the observed $t$-statistic $x^{\text{obs}}$ and then inverting $\theta\mapsto F(x^{obs};\theta,\bar{x})$.
Our main result is for situations where $x^{\text{obs}}$ only barely exceeds $\bar{x}$, a case often referred to as “marginal significance”. For this case, we show that the conditional confidence interval is very wide and located far away from the estimated value of the parameter. Specifically, we show that even the upper bound of the interval diverges to $\infty$- ∞ as $x^{\text{obs}}\downarrow\bar{x}$. Therefore, at the margin of significance, the confidence interval contains only extremely negative numbers, a range of numbers that may be a-priori implausible.
Our results are closely related to Kivaranovic & Leeb (2021), who show that the expected length of a confidence interval constructed by the polyhedral method is infinite when the conditioning set is bounded either from below or from above. Indeed, our main result immediately follows from their Lemma A.3. However, our focus differs from theirs. They consider the expected length of the confidence interval. We complement their results by identifying precisely when (that is, in what kind of samples) the confidence interval exhibits extreme behavior. Whereas Kivaranovic & Leeb (2021) focus on the length of the confidence interval, our focus is on its location. We show that the conditional confidence interval for a marginally significant parameter is located far away from its estimate. This result applies more broadly than our file drawer problem, since it builds on Lemma A.3 of Kivaranovic & Leeb (2021) which holds generally for truncated normal distributions. In particular, when the conditioning set is bounded from either below or above and the observed parameter is at the margin of the conditioning set, the conditional selective confidence intervals are located in extreme ranges.
As far as we know, this “location problem” of selection-robust inference on marginally significant parameters has not been previously discussed in the literature. It is of great practical importance, given that empirical $t$-statistics often cluster close to common critical values, as documented by Gerber & Malhotra (2008), Brodeur et al. (2016), and Brodeur et al. (2020). This clustering has raised concerns about potential $p$-hacking and led to marginally significant results being viewed with skepticism. Selection-robust inference methods may be able to address some of the concerns about marginally significant results. However, the “location problem” implies that the difference between robust estimates and the original (conventional) estimates may be extreme.
A parameter estimate can be considered “highly significant” when its $t$-statistic far exceeds the critical value. In this case, the conditional selective confidence interval is almost identical to the conventional interval. This feature has already been documented in the literature. For instance, Andrews et al. (2024) discuss statistical inference for determining the best welfare-improving policy and show that the confidence interval based on the polyhedral method is nearly identical to the conventional one when the best policy is unambiguously determined.
For moderate cases, where the parameter estimate is neither “marginally” nor “highly” significant, the conditional selective confidence interval provides non-trivial insights. It is wider than the conventional interval, but still informative enough to provide a useful range of parameter values. The increased length intuitively reflects additional statistical uncertainty when the data it is used both for selection and inference.
We show that the problem of extreme confidence intervals at the “margin of significance” does not appear when statistical significance is determined by a two-sided test, that is, a test that rejects if $\lvert x^{\text{obs}}\rvert>\bar{x}$. In this case, the limiting interval as $x^{\text{obs}}\downarrow\bar{x}$ is located around zero and has a relatively short length. This is already suggested by the result of Kivaranovic & Leeb (2021) that the expected length of the interval is finite in this case. A perhaps surprising result of our analysis is that, even though the interval is “well-behaved” in the sense that it does not diverge, it may still exclude the conventional estimate of the parameter.
Our arguments are entirely mechanical and do not rely on the underlying distribution of the data. Regardless of the data’s true distribution or the parameter’s actual value, if the observed value of the $t$-statistic is close to the critical value, the problem identified in this paper will arise. This aspect of our results starkly contrasts other studies that highlight potential issues with selective inference methods. For instance, the result in Kivaranovic & Leeb (2021) on the expected length relies on distributional assumptions to evaluate the expectation. Our result does not need any distributional assumptions.
The remainder of the paper is organized as follows. Section 1 introduces the settings. The problems are illustrated via a numerical example in Section 2. Section 3 presents the main analytical results. We discuss the two-sided significance cases in Section 4. Section 5 concludes.
1 Settings
We consider the file drawer problem from Rosenthal (1979) of reporting only statistically significant effects. The selective conditional inference procedure based on the polyhedral approach Lee et al. (2016) is introduced.
We are interested in a parameter $\theta$ for which an estimator $\hat{\theta}$ is available. Assume that the estimator $\hat{\theta}$ is unbiased and Gaussian: $\hat{\theta}\sim N(\theta,\sigma^{2})$. For simplicity, we assume that the variance $\sigma^{2}$ is known. In this case, it is without loss of generality to normalize $\sigma^{2}=1$ and $\hat{\theta}$ is identical to its $t$-statistic $X=X/\sigma$. Let $x^{\text{obs}}$ denote the observed value of $X$. First, we discuss the one-sided significance, where an effect is significant if $x^{\text{obs}}>\bar{x}$ for a pre-specified critical value $\bar{x}$. The two-sided significance case is discussed in Section 4.
We follow the polyhedral approach given in Lee et al. (2016). It is based on the conditional distribution of $X$ given that it exceeds the critical value. Under the Gaussianity assumption, the conditional distribution function is given by
$F(a;\theta,\bar{x}):=$ $\Pr(X\leq a\mid X\geq\bar{x})$
$=$ $\frac{\Phi(a-\theta)-\Phi(\bar{x}-\theta)}{1-\Phi(\bar{x}-\theta)}.$
Let $\theta(p)$ denote the value of $\theta$ that solves the equation
$p=F(x^{\text{obs}};\theta,\bar{x}),$ (1)
for given $p\in(0,1)$, $x^{\text{obs}}$, and $\bar{x}$. The solution exists by Lemma A.2 in Kivaranovic & Leeb (2021) and is strictly decreasing in $p$. The dependence of $\theta(p)$ on $x^{\text{obs}}$ is crucial for the results in this paper, even though it is kept implicit in the notation.
Various statistics for the selective conditional inference correspond to $\theta(p)$ with specific choices of $p$. The conditionally median unbiased estimator is given by
$\hat{\theta}_{MU}:=\theta(0.5).$
A $100(1-\alpha)$conditional confidence interval is defined as
$[\theta(1-\alpha/2),\theta(\alpha/2)].$
For example, $[\theta(0.975),\theta(0.025)]$ is the 95% confidence interval.
To verify that the confidence interval has the desired conditional coverage probability, note that
$\Pr(\theta\in[\theta(1-\alpha/2),\theta(\alpha/2)]\mid X\geq\bar{x})$
$=$ $\Pr(\alpha/2\leq F(X,\theta,\bar{x})\leq 1-\alpha/2\mid X\geq\bar{x})=1-\alpha,$
where the last equality holds because the Rosenblatt (1952) transformation implies that, conditional on $X\geq\bar{x}$, $F(X,\theta,\bar{x})$ is uniformly distributed on the unit interval.
2 A numerical illustration
We now present a numerical illustration of the issues discussed in this note for the case of one-sided significance testing. The critical value is set to $\bar{x}=1.64$, corresponding a nominal level of 5%. We set the targeted (conditional) coverage of the confidence intervals to 95%.
Figure 1: Test statistic and confidence intervals
NOTE: Conditional inference after one-sided testing with critical value $\bar{x}=1.64$. The figure plots the median-unbiased estimator (solid curve) and the bounds $\theta(0.975)$ and $\theta(0.025)$ of the 95% conditional confidence interval (bold dashed curves). The thin dashed line is the 45-degree line, representing the unconditional estimate. The thin dash-dotted lines give the usual confidence interval.
Figure 1 plots the median-unbiased estimator and the conditional confidence interval as a function of the observed $t$-statistic $x^{\text{obs}}$. The upper dashed curve represents the upper bound $\theta(0.025)$ of the confidence interval and the lower dashed curve represents the lower bound $\theta(0.975)$. The solid curve represents the median unbiased estimator $\hat{\theta}_{MU}$. The dashed line is the 45-degree line. Since we have normalized $X=\hat{\theta}$, it gives the unconditional estimate $\hat{\theta}$. The two parallel dash-dotted lines represent the conventional confidence interval $x^{\text{obs}}\pm 1.96$.
Figure 1 reveals several noteworthy insights. When the estimated parameter is highly significant, i.e., when $x^{\text{obs}}$ is large, selective inference yields almost identical results to conventional inference that ignores the selection problem. In this case, the median-unbiased estimator is close to the conventional parameter estimate ($\hat{\theta}_{MU}\approx\hat{\theta}$) and the conditional confidence interval approximates the traditional equal-tailed interval ($\theta(0.975)\approx x^{\text{obs}}-1.96$ and $\theta(0.025)\approx x^{\text{obs}}+1.96$).
When the observed $t$-statistic is of moderate size, conditional inference results differ from the conventional inference. The median-unbiased estimator is lower than the conventional estimate. The conditional confidence interval is wider than the conventional one. Its lower bound is below the lower bound of the standard interval, whereas the upper bounds of the two intervals are fairly similar. The downward adjustments reflect correction for selection bias. Significant effects are positively selected since overestimated effects are more likely to clear the threshold of significance. The wider interval reflects an increase in statistical uncertainty when the data is used for both parameter selection and inference.
In the case of marginal significance, that is, when $x^{\text{obs}}$ is close to the threshold $\bar{x}$, the downward corrections are quite extreme. As $x^{\text{obs}}$ approaches $\bar{x}$, the upper bound of the conditional confidence interval drops below the critical value $\bar{x}$ and eventually below zero. This property of conditional inference may seem counterintuitive. Significantly positive parameter estimates are often interpreted as providing a “statistical guarantee” of a positive population parameter. This perspective seems incompatible with inferring a negative range of parameter values. However, the guarantees given by significance testing are unconditional and do not contradict the inference results based on the conditional distribution. Still, from a practical perspective, a confidence interval that gives a range of negative values may be deemed implausible and one can simply state that the inference results are uninformative.
The numerical results in Figure 1 are easily analytically verified. For example, the median-unbiased estimator is smaller than the critical value $\bar{x}$ if $\theta(0.5)<\bar{x}$ and the upper-bound of the confidence interval is smaller than $\bar{x}$ if $\theta(0.025)<\bar{x}$. The solution $\theta(p)$ depends on the observed $t$-statistic $x^{\text{obs}}$. To find the range of $x^{\text{obs}}$ such that $\theta(p)<\bar{x}$, solve the inequality
$p>F(x^{\text{obs}};\bar{x},\bar{x})=\frac{\Phi(x^{\text{obs}}-\bar{x})-0.5}{0.5},$
for $x^{\text{obs}}$ to obtain
$x^{\text{obs}}<\Phi^{-1}(0.5p+0.5)+\bar{x}.$
Thus, unless $x^{\text{obs}}$ exceeds $\bar{x}$ by more than $\Phi^{-1}(0.5p+0.5)$, $\theta(p)$ is smaller than $\bar{x}$. For example, if $x^{obs}<\Phi^{-1}(0.5\times 0.025+0.5)+\bar{x}\approx 0.031+\bar{x}$, the upper bound of the 95% confidence interval is less than $\bar{x}$.
Similarly, we can show that we have $\theta(p)<0$ for
$x^{\text{obs}}<\Phi^{-1}((1-\Phi(\bar{x}))p+\Phi(\bar{x})).$
Thus, unless $x^{\text{obs}}$ exceeds $\Phi^{-1}((1-\Phi(\bar{x}))p+\Phi(\bar{x}))$, $\theta(p)$ is negative. In particular, the upper bound of a 95% confidence interval is negative when $x^{\text{obs}}<\Phi^{-1}(0.05\times 0.025+0.95)\approx 1.66$. This means that the estimate $\hat{\theta}$ is significantly positive at the selection stage, but significantly negative at the inference stage. This apparent contradiction can be reconciled by noting that the selection stage is based on the unconditional distribution of $\hat{\theta}$, whereas the inference stage is based on its conditional distribution.
3 Theoretical results
In this section, we analytically derive the behavior of conditional selective inference under the “marginally significant” and “highly significant” cases. For the former, we consider the behavior of $\theta(p)$ for large values of the observed $t$-statistic $x^{\text{obs}}$. For the latter, we consider the behavior of $\theta(p)$ when $x^{\text{obs}}$ is close to the critical value $\bar{x}$.
We first consider the case of marginal significance. We show that the bounds of the conditional selective confidence intervals and the median-unbiased estimator can take arbitrarily large negative values when $x^{\text{obs}}$ is close to $\bar{x}$. The following theorem is a direct application of Lemma A.3 in Kivaranovic & Leeb (2021).
Theorem 3.1.
For $p\in(0,1)$,
$\theta(p)\to-\infty,$
as $x^{\text{obs}}\downarrow\bar{x}$,
Proof.
See Lemma A.3 in Kivaranovic & Leeb (2021). ∎
This result holds regardless of the true distribution of $X$. Its proof is entirely based on the functional form of $F(a,\theta,\bar{x})$ and does not rely on the probabilistic characteristics of the sample.
Theorem 3.1 implies that the upper bound of the conditional confidence approaches negative infinity as the estimated effect size nears the point of insignificance. Specifically, at the margin of significance, the confidence interval contains only extremely large negative values. In many practical applications, such extreme values are a-priori implausible.
Turning to the case of a highly significant parameter estimate, note that the conventional equal-tailed confidence interval at the $(1-\alpha)$-level is given by the interval $[\theta^{*}(1-\alpha/2),\theta^{*}(\alpha/2)]$ with $\theta^{*}(p)=x^{\text{obs}}-\Phi^{-1}(p)$. For example, the conventional 95% confidence interval is $[x^{\text{obs}}-1.96,x^{\text{obs}}+1.96]$. Similarly to $\theta(p)$, $\theta^{*}(p)$ can also be found by inverting a distribution function. In particular, $\theta^{*}(p)$ is the value of $\theta$ that solves the equation $p=\Phi(x^{\text{obs}}-\theta)$. The conventional point estimate is given by $\hat{\theta}=\theta^{*}(0.5)$.
The following result establishes that, as $x^{\text{obs}}\uparrow\infty$, the conditional selective confidence interval converges to the conventional confidence interval and the conditional median-unbiased estimate converges to the conventional estimate.
Theorem 3.2.
For $p\in(0,1)$,
$\theta(p)-\theta^{*}(p)\to 0$
as $x^{\text{obs}}\uparrow\infty$.
Proof.
Fix $p\in(0,1)$ and take $\xi>0$ arbitrarily. To show that $\lvert\theta(p)-\theta^{*}(p)\rvert\leq\xi$ eventually, let $\delta>0$ be small enough that
$\Phi^{-1}(p+\delta)-\Phi^{-1}(p-\delta)\leq\xi.$
We can find such $\delta$ by the continuity of $\Phi^{-1}$. Since $\theta^{*}(a)=x^{\text{obs}}-\Phi^{-1}(a)$, it suffices to show that, for $x^{\text{obs}}$ large enough,
$\theta^{*}(p+\delta)\leq\theta(p)\leq\theta^{*}(p-\delta).$
Because $\bar{x}$ is fixed, for every $\epsilon>0$, there exists $M(\epsilon)$ such that for any $x^{\text{obs}}>M(\epsilon)$,
$\Phi(\bar{x}-(x^{\text{obs}}-\Phi^{-1}(p)))<\epsilon.$ (2)
Let $\epsilon>0$ be small enough that
$(p-\delta)/(1-\epsilon)<p<p+\delta-\epsilon.$
By inequality (2), for $x^{\text{obs}}>M(\epsilon)$,
$F(x^{\text{obs}};\theta^{*}(p-\delta),\bar{x})\leq$ $\frac{\Phi(x^{\text{obs}}-\theta^{*}(p-\delta))}{1-\epsilon}$
$\leq$ $\frac{p-\delta}{1-\epsilon}<p.$
Since $F(x^{\text{obs}};\theta(p),\bar{x})=p$ and $F(x^{\text{obs}};\theta,\bar{x})$ is strictly decreasing in $\theta$ by Lemma A.2 in Kivaranovic & Leeb (2021), this establishes $\theta(p)\leq\theta^{*}(p-\delta)$ for $x^{\text{obs}}>M(\epsilon)$. Also, by inequality (2), for $x^{\text{obs}}>M(\epsilon)$,
$F(x^{\text{obs}};\theta^{*}(p+\delta),\bar{x})\geq$ $\Phi(x^{\text{obs}}-\theta^{*}(p+\delta))-\epsilon$
$=$ $p+\delta-\epsilon>p$
and therefore $\theta(p)\geq\theta^{*}(p+\delta)$ for $x^{\text{obs}}>M(\epsilon)$, concluding the proof. ∎
4 Two-sided significance test
In the case of a two-sided test, the conditional selective confidence interval is not located in an extreme range. We consider the setting in Example 1 of Fithian et al. (2017) where the estimate $\hat{\theta}$ is reported only when $|x^{\text{obs}}|\geq\bar{x}$ for a critical value $\bar{x}$.
Conditional inference can be conducted similarly to the one-sided case with a different conditional distribution function. Conditionally on being reported, the distribution function of $\hat{\theta}=X$ is given by
$F_{\text{2-sided}}(a,\theta,\bar{x})=\Pr(X\leq a\mid|X|\geq\bar{x})$
$=$ $\frac{\min(\Phi(a-\theta),\Phi(-\bar{x}-\theta))+\mathbf{1}_{a>\bar{x}}(\Phi(a-\theta)-\Phi(\bar{x}-\theta))}{\Phi(-\bar{x}-\theta)+\Phi(-\bar{x}+\theta)},$
where $\mathbf{1}_{a>\bar{x}}=1$ if $a>\bar{x}$ and $=0$ otherwise. Let $\theta_{\text{2-sided}}(p)$ denote the value of $\theta$ that solves $p=F_{\text{2-sided}}(x^{\text{obs}},\theta,\bar{x})$.
The median-unbiased estimator and conditional confidence interval are defined analogously to the one-sided significance case. In particular, the median-unbiased estimator is given by $\hat{\theta}_{MU,2}=\theta_{\text{2-sided}}(0.5)$ and a $(1-\alpha)$-level confidence interval is given by $[\theta_{\text{2-sided}}(1-\alpha/2),\theta_{\text{2-sided}}(\alpha/2)]$.
In contrast to the one-sided significance case, $\theta_{\text{2-sided}}(p)$ converges to a finite limit as $x^{\text{obs}}\downarrow\bar{x}$. For example, at $p=0.5$, the limit solves the equation $0.5=\Phi(-\bar{x}-\theta)/(\Phi(-\bar{x}-\theta)+\Phi(-\bar{x}+\theta))$ and the median-unbiased estimate is given by $\hat{\theta}_{MU,2}=0$ as $x^{\text{obs}}\downarrow\bar{x}$.
Figure 2: Test statistic and confidence intervals conditional on two-sided significance
NOTE: Conditional inference after two-sided testing with critical value $\bar{x}=1.64$. The figure plots the median-unbiased estimator (solid curve) and the bounds $\theta(0.975)$ and $\theta(0.025)$ of the 95% conditional confidence interval (bold dashed curves). The thin dashed line is the 45-degree line, representing the unconditional estimate. The thin dash-dotted lines give the usual confidence interval.
Figure 2 displays the conditionally median-unbiased estimator and the upper and lower bounds of a 95% conditional selective confidence interval. The range of the horizontal axis are positive values of $x^{\text{obs}}$ that lead to a significant parameter estimate. The corresponding graphs for $x^{\text{obs}}<-\bar{x}$ are obtained by mirroring the figure around the vertical axis. The figure demonstrates several key features of conditional inference after selection based on two-sided significance.
First, the median-unbiased estimator is close to zero when the estimate is marginally significant. This stands in contrast to the one-sided significance case, where the median-unbiased estimator diverges to negative infinity.
Secondly, in the case of a highly significant parameter estimate, the conditional selective confidence interval is similar to the conventional confidence interval. This is the same result as in the one-sided significance case.
Thirdly, the lower bound of the conditional confidence interval is below the lower bound of the conventional confidence interval but not to an extreme extent. The largest difference in the lower bounds is about $0.847$, occurring at $x^{\text{obs}}\approx 2.8$.
Fourthly, the upper bound of the 95% conditional selective confidence interval falls below the conventional upper bound when $x^{\text{obs}}$ is small (for example, when it is smaller than 2.0). Very close to the critical value $\bar{x}$, the upper bound falls even below the estimate $\hat{\theta}=x^{\text{obs}}$. At this point, the conventional estimate is excluded from the conditional selective confidence interval.
Lastly, the length of the conditional selective confidence interval can be narrower than the conventional interval. Given that the upper bound is at most slightly above the conventional upper bound, the length of the conditional confidence interval is bounded and does not vastly exceed the length of the conventional confidence interval. As $x^{\text{obs}}$ approaches $\bar{x}$ from above, the length of the confidence interval decreases. The largest difference between the conditional and conventional intervals is about $2.15$ and occurs around the threshold $x^{\text{obs}}\approx\bar{x}=1.64$ where the conditional interval is the narrowest. Here, the length of the unconditional confidence interval is $3.92(=1.96\times 2)$, and that of the conditional one is $1.77$; the difference is substantial.
5 Conclusion
In this paper, we explore the properties of selective inference methods. Our main finding is that the results of these inference methods can be extreme when the parameter is only marginally significant. Therefore, applied researchers should use selective inference methods with caution. These methods should only be employed when the test statistic sufficiently exceeds the critical value.
While we consider a simple file drawer problem in this paper, results analogous to that presented in Section 3 can be obtained for general polyhedral methods, provided that the conditioning set is bounded either from below or from above. This conclusion is drawn from Lemma A.3 of Kivaranovic & Leeb (2021), which applies to general polyhedral methods. For example, conditional inference after variable selection by Lasso is affected by this issue.
In contrast, when the conditioning set is bounded, as in the case of inferences after two-sided significance testing, the problem is less pronounced. While we do not study other conditional selective inference methods such as data carving (Fithian et al., 2017) and selection with a randomized response (Tian & Taylor, 2018) in this paper, these methods yield confidence intervals whose expected lengths are bounded (Kivaranovic & Leeb, 2024). To the best of our knowledge, the location characteristics of their confidence intervals have not yet been investigated in the literature. We conjecture that the problem manifests less dramatically with these methods.